OpenStudy (anonymous):

Where is the y-intercept for the line 3x + 2y + 8 = 0 A 2,0 B 0,8 C 0, -4 D -4, 0

Vocaloid (vocaloid):

when x = 0, what does y = ?

OpenStudy (anonymous):

Turn the equation into the y-intercept form. y-intercept form y = mx +b m = slope b = y-intercept \( \huge 3x + 2y + 8 = 0 \) \( \huge 3x + 8 -8 = -8 \) \( \huge 3x + 2y = -8 \) \( \huge 3x - 3x + 2y = -3x - 8 \) \( \huge 2y = -3x -8 \) \( \huge \frac{2y}{ 2} = \frac{-3x}{2} -\frac{8}{ 2} \) \( \huge y = -\frac{3x}{2} - 4 \) Now this is in y-intercept form y = mx + b. b is your y value to the y-intercept form. Remember y-intercept = the point (0, y) so what is the y intercept point. You can also find the y-intercept by input 0 into the equation like @Vocaloid stated. suggested.